Please help me with this question, this is for an upcoming test I'm preparing for. I really need help! Thank you so much!
Question:
Find two whole numbers whose product is 147 and whose quotient is 3.
I tried to do the question by myself and got: "The two whole numbers are 7 and 21.
I'm not sure if the answer is correct but I tried my best. However, I still need to find out how to get to the answer and all the work and steps involved in getting the answer. Please show me a formula or the procedure tot get to the answer. Thank you so much, I really appreciate it! :D

Answer by stanbon(75887)

Find two whole numbers whose product is 147 and whose quotient is 3.
-----------------
Equations:
Product: xy = 147
x/y = 3
---
Solve the 2nd for "x":
x = 3y
----
Substitute into the 1st to get:
(3y)y = 147
3y^2 = 147
y^2 = 49
y = 7
-----
Solve for "x":
x/y = 3
x/7 = 3
x = 21
---------
Solution:
The numbers are 7 and 21
==============================
Cheers,
Stan H.
===============